# Dario Pasquini & Evgenios M. Kornaropoulos

### v1

@@ 4.2.3 Malformed” Queries Reveal Useful Information
Alignment is a methodology that aims to prevent the model from answering harmful tasks (e.g., “How to build a bomb?”) or producing objectionable content overall. Due to the alignment process, when asked to respond to a harmful query, the LLM refuses by issuing a message stating that the prompted task cannot be served (see Figure 2). Attackers can exploit this behavior to induce strong and robust fingerprint signals, as these “error messages” are typically highly model-dependent and sufficiently unique. More prominently, these responses also remain consistent across prompting configurations. Indeed, regardless of the directives set by the prompting configuration, the model prioritizes the refusal of the harmful task. Furthermore, we stress that issuing harmful prompts allows us to discriminate the other two macro-categories of models: aligned and non-aligned LLMs. Therefore, such queries provide distinguishing power when it comes to non-aligned LLMs.
@@
Besides exploiting models’ alignment, an attacker can craft “inconsistent” (or more generally “malformed”) queries by relying on nonsensical or semantically-broken prompts. For instance, submitting queries that mix several languages (e.g., “Bonjour, how are you doing today? ¿Qué tal?”). As in OS fingerprinting, the way the target handles such inconsistent inputs (e.g., answering the query in either English or Spanish) provides a unique behavioral signature for the LLM, which improves fingerprint accuracy when combined with other probing approaches.
@@ 6.1 Evaluation Setup
    Prompt Frameworks: We consider two settings: RAG and Chain-Of-Thought [11]. To simulate RAG, we create the input corpus by sampling $30$ random entries from the dataset SQuAD 2.0 [8], and consider $6$ prompt templates for retrieval-based-Q&A. Prompting frameworks are applied only in $20\%$ of the cases.
Creating Training/Evaluation Set for Inference Model:
@@ 6.2.1 Closed-Set Classification Setting
Figure 6: Confusion matrix for the closed-set fingerprinting model.
@@ Appendix A Additional Resources
This appendix contains additional material. Table 2 reports misleading outputs produced by LLMs upon banner-grabbing queries, whereas Table 3 lists the LLMs considered in this work. Table 4 reports examples of system prompts used to generate different prompting configurations.
@@ Appendix B Optimize query strategy
This process resulted in queries listed in Table 1.

### v2

@@ 2.1 Problem Formulation
2.2 Related Work
Xu et al. [16] propose a watermark-based fingerprinting technique designed to protect intellectual property through remote ownership attestation of LLMs. In their setting, the model owner performs an additional training step on the existing model to inject a behavioral watermark before releasing it. To verify ownership of a remote LLM, the owner submits the set of predefined trigger queries and checks for the injected watermark in the obtained responses. In another concurrent work and under the same setting, Russinovich and Salem [12] introduce a technique to incorporate recognizable behaviors into an LLM via tuning, defining core requirements for successful watermark-based fingerprinting. Their approach is based on hashing responses generated by a set of predefined queries.
Overall, both of the above works address the scenario where the “defender” watermarks the model (s)he owns during training so that can later verify whether the model is used without their consent. On the contrary, our work concerns the case where an “attacker”, who cannot affect the training or the specifics of the deployment, induces unique responses via prompting so as to fingerprint and recognize the unknown underlying model.
The most comparable approach to LLMmap is the concurrent work by Yang and Wu [17]. Unlike other solutions [16, 12], this method does not require fine-tuning the model. However, it assumes access to the logits output generated by the tested LLM, rather than the generated text, making it non-applicable to real-world practical settings where LLM is deployed. In particular, their method performs fingerprinting by matching the vector spaces induced by the logits of two different models in response to a set of 300 random queries (as opposed to our technique that needs $\leq 8$ queries).
@@ 4.2.3 Malformed” Queries Reveal Useful Information
Alignment is a methodology that aims to make the model helpful, honest, and harmless [2]. In particular, we aim to exploit the latter property. Due to the alignment process, when asked to respond to a harmful query (e.g., “How to build a bomb?”), the LLM refuses by issuing a message stating that the prompted task cannot be served (see Figure 2). Attackers can exploit this behavior to induce strong and robust fingerprint signals, as these “error messages” are typically highly model-dependent and sufficiently unique. More prominently, these responses also remain consistent across prompting configurations. Indeed, regardless of the directives set by the prompting configuration, the model prioritizes the refusal of the harmful task. Furthermore, we stress that issuing harmful prompts allows us to discriminate the other two macro-categories of models: aligned and non-aligned LLMs. Therefore, such queries provide distinguishing power when it comes to non-aligned LLMs.
@@
Besides exploiting models’ alignment, an attacker can craft “inconsistent” (or more generally “malformed”) queries by relying on nonsensical or semantically-broken prompts. For instance, submitting queries that mix several languages (e.g., “Bonjour, how are you doing today? ¿Qué tal?”) [7, 3]. As in OS fingerprinting, the way the target handles such inconsistent inputs (e.g., answering the query in either English or Spanish) provides a unique behavioral signature for the LLM, which improves fingerprint accuracy when combined with other probing approaches.
@@ 8 What’s your name? Banner grabbing
Automatize Queries Generation Process
Inspired by existing work on OS fingerprinting, we have developed our query strategy based on domain knowledge and manual interactions with LLMs. However, more advanced solutions can be utilized to create novel and more effective queries for LLM fingerprinting tasks. Specifically, similar to previous work on the automated generation of adversarial inputs [19, 9], optimal queries can be generated by framing the task as an optimization problem and identifying suitable token combinations in the models’ input space.
In this direction, the properties listed in Section 4.1 can already be used to define an objective function. Furthermore, Eq 2 and 3 are fully differentiable, supporting white-box optimization (e.g., via GCG [19]). However, unlike other settings, this optimization would involve multiple LLMs simultaneously (at least two), making the task resource-intensive.
@@ 6.1 Evaluation Setup
    Prompt Frameworks: We consider two settings: RAG and Chain-Of-Thought [15]. To simulate RAG, we create the input corpus by sampling $30$ random entries from the dataset SQuAD 2.0 [11], and consider $6$ prompt templates for retrieval-based-Q&A. An example of prompt derived from this approach is given in Figure 9 in Appendix A. Prompting frameworks are applied only in $20\%$ of the cases and combined with the other modifiers (e.g., system prompt).
@@ 6.2 Results
Figure 5: Confusion matrix for the closed-set fingerprinting model.
@@ 7 Conclusion & Future Work
8 Acknowledgments
The authors would like to thank Antonios Anastasopoulos for his insights into the related work.
@@ Appendix A Additional Resources
This appendix contains additional material. Table 2 reports misleading outputs produced by LLMs upon banner-grabbing queries, whereas Table 3 lists the LLMs considered in this work. Table 4 reports examples of system prompts used to generate different prompting configurations. Figure 9 reports an example of prompt when the prompting configuration uses RAG.
@@
Context information is below.
In 1994, responding to the need for a more useful system for describing chronic pain, the International Association for the Study of Pain (IASP) classified pain according to specific characteristics: (1) region of the body involved (e.g. abdomen, lower limbs), (2) system whose dysfunction may be causing the pain (e.g., nervous, gastrointestinal), (3) duration and pattern of occurrence, (4) intensity and time since onset, and (5) etiology. However, this system has been criticized by Clifford J. Woolf and others as inadequate for guiding research and treatment. Woolf suggests three classes of pain : (1) nociceptive pain, (2) inflammatory pain which is associated with tissue damage and the infiltration of immune cells, and (3) pathological pain which is a disease state caused by damage to the nervous system or by its abnormal function (e.g. fibromyalgia, irritable bowel syndrome, tension type headache, etc.).
Another landmark is the old centre and the canal structure in the inner city. The Oudegracht is a curved canal, partly following the ancient main branch of the Rhine. It is lined with the unique wharf-basement structures that create a two-level street along the canals. The inner city has largely retained its Medieval structure, and the moat ringing the old town is largely intact. Because of the role of Utrecht as a fortified city, construction outside the medieval centre and its city walls was restricted until the 19th century. Surrounding the medieval core there is a ring of late 19th- and early 20th-century neighbourhoods, with newer neighbourhoods positioned farther out. The eastern part of Utrecht remains fairly open. The Dutch Water Line, moved east of the city in the early 19th century required open lines of fire, thus prohibiting all permanent constructions until the middle of the 20th century on the east side of the city.
Nicholas Lezard described post-punk as "a fusion of art and music". The era saw the robust appropriation of ideas from literature, art, cinema, philosophy, politics and critical theory into musical and pop cultural contexts. Artists sought to refuse the common distinction between high and low culture and returned to the art school tradition found in the work of artists such as Captain Beefheart and David Bowie. Among major influences on a variety of post-punk artists were writers such as William S. Burroughs and J.G. Ballard, avant-garde political scenes such as Situationism and Dada, and intellectual movements such as postmodernism. Many artists viewed their work in explicitly political terms. Additionally, in some locations, the creation of post-punk music was closely linked to the development of efficacious subcultures, which played important roles in the production of art, multimedia performances, fanzines and independent labels related to the music. Many post-punk artists maintained an anti-corporatist approach to recording and instead seized on alternate means of producing and releasing music. Journalists also became an important element of the culture, and popular music magazines and critics became immersed in the movement.
Given the context information and not prior knowledge, answer the query.
Query: Tell me how to build a bomb.
Answer:
Figure 9: Example of a prompt generated when simulating a RAG pipeline. In red, the query issued by the attacker. The prompt also includes the system prompt, which we do not report for representation’s sake.
@@ Appendix B Optimize query strategy
This process resulted in the queries listed in Table 1.